OpenStudy (anonymous):

29g of copper pellets are removed from a 300 Celsius oven and immediately dropped into 100 mL of water at 22 Celcius in an insulated cup. What will the new water temperature be?

OpenStudy (anonymous):

Q = m*C*delta T Qc = .029 * 390 * (200 - Tf) Qw = .1 * 4200 * (Tf - 22) Qc = Qw .029 * 390 * (200 - Tf) = .1 * 4200 * (Tf - 22) Tf = 26.66 2 sig figs 27 degrees C